# From Cash to Cashless: UPI’s Impact on Spending Behavior among Indian Users

### v2

@@ 4.2.4. RECOMMENDATIONS FOR FINANCE TRACKING APPLICATIONS
5. Discussion
Answering RQ1 as mentioned in Section 1, the study results reveal a significant impact of UPI on Indian users’ spending habits. Most respondents (74.2%) accepted that UPI adoption led to increased spending, highlighting the UPI influence on user behavior. The data also indicates that many respondents started using UPI during significant economic and social changes, such as demonetization and pandemic-related lockdowns. Furthermore, 59.8% of respondents acknowledged exceeding their budget or overspending due to UPI usage. The qualitative analysis supports these findings, indicating that UPI has contributed to increased spending for most users. The abstract nature of digital transactions leads to psychological shifts, contributing to impulsive spending. In some cases, regret and guilt hit people after they realize their spending by checking their bank balances or statements, compared to spending physical cash, highlighting the lack of tangibility associated with UPI payments.
5.1. RECOMMENDATIONS AND DESIRED FEATURES IN UPI & FINANCE TRACKING APPLICATIONS
Our research insights highlight key recommendations for UPI platforms and financial app developers, aiming to elevate the digital payment experience. Users increasingly seek budgeting and expenditure limitation features within UPI apps, enabling them to set spending thresholds for specific categories and receive timely reminders to stay within financial goals. Visualization tools like graphical representations and pie charts should be integrated to enhance user comprehension, offering a clear overview of spending patterns over time.
In addition, there is a growing demand for educational content on responsible financial management embedded within the apps to empower users with financial literacy. Recognizing the diverse needs of different professions is crucial; for instance, business professionals may require tools to manage client payments and categorize expenses efficiently. Tailoring UPI solutions to cater to various user segments ensures accessibility and relevance.
5.2. FINANCIAL APP ANALYSIS IN INDIA
@@
Our comprehensive analysis of various financial tracking and budget management apps in the Indian market is summarized in Table 1. These applications offer diverse features tailored to meet individual financial requirements, ranging from automated organization of transactional data to robust expense-tracking capabilities. Users can leverage these tools to elevate their financial management skills. However, some notable gaps exist in meeting the recommendations outlined in Section 5.1. Specific features like detailed graphical representations of expenditure patterns and spending thresholds with reminders are unevenly available across platforms. Providing educational content on responsible financial management also varies, with certain apps lacking sufficient resources. Addressing these gaps requires finer control mechanisms for spending limits, real-time notifications, goal-based savings integration, AI-powered financial insights, customizable reporting, collaborative budgeting, integration with financial education platforms, and improvements in expense categorization. Enhancing user convenience should involve autonomously pulling expense data without manual input, ensuring a seamless experience while safeguarding user privacy. Balancing automation with privacy protection is crucial for advancing a comprehensive and empowering financial tracking solution for users.
@@
In conclusion, while existing financial tracking apps offer valuable features, there is room for improvement in aligning with the recommendations outlined. Future developments should focus on bridging these gaps to provide users with a more holistic and empowering financial management experience.
Stakeholders, including app developers, government agencies, and users, stand to benefit from incorporating our suggestions into future financial tracking app developments. This comprehensive analysis addresses user needs and provides valuable insights for enhancing the financial management ecosystem, ultimately advancing effective and user-centric solutions.

### v3

@@ 5.3.4. Recommendations for Finance Tracking Applications
6. Insights from Phase 1 Evaluation
Users expressed a desire to incorporate customizable spending limits and real-time notifications into the redesigned app, ensuring users stay aligned with their financial goals. In addition, users frequently requested better ways to visualize their spending patterns with the integration of detailed graphical tools, including pie charts, providing users with a clear representation of their financial habits over time. Financial education also emerged as a significant gap in many apps, with users seeking educational pop-ups and financial tips with a built-in “Financial Education” component which offers users practical financial quotes and suggestions through timely pop-ups during transactions, helping them stay informed on budgeting, saving, and responsible spending. By delivering these insights in real-time, the app can enhance users’ financial literacy and promote better decision-making. Lastly, our results highlighted a strong user preference for automated expense tracking that balances privacy concerns associated with accessing the financial details of the users.
@@ 6. Insights from Phase 1 Evaluation
Our comprehensive analysis of various financial tracking and budget management apps in the Indian market is summarized in Table 1. These applications offer diverse features tailored to meet individual financial requirements, ranging from automated organization of transactional data to robust expense-tracking capabilities. Users can leverage these tools to elevate their financial management skills. Specific features like detailed graphical representations of expenditure patterns and spending thresholds with reminders are unevenly available across platforms.
@@
We examined which features already exist in these apps, particularly those related to budget tracking, spending categorization, and financial planning. A key observation was that although many of these apps excel in budget tracking and expense management, none of them support UPI payments. This presents a significant limitation, as users are forced to switch between multiple apps to manage expenses and make payments. In conclusion, while existing financial tracking apps offer valuable features, there is room for improvement in aligning with the recommendations outlined. Future developments should focus on bridging these gaps to provide users with a more holistic and empowering financial management experience.